OpenStudy (anonymous):

if possibl,m ealuatef(x) at the given values of x f(x0=x^1/4 x=16,-10

OpenStudy (anonymous):

\[f(x)=x^{\frac{1}{4}}=\sqrt[4]{x}\]

OpenStudy (anonymous):

so \[f(16)=\sqrt[4]{16}=2\]

OpenStudy (anonymous):

and \[f(-10)\] does not exist

OpenStudy (anonymous):

thank u.....